OpenStudy (anonymous):

What is the vertex of the graph y=x^2-11x+28

OpenStudy (anonymous):

use \[-\frac{b}{2a}\] with \[a=1,b=-11\] to get \[\frac{11}{2}\]

OpenStudy (anonymous):

that is first coordinate. second is what you get when you replace x by 11/2

OpenStudy (anonymous):

i get \[-\frac{9}{4}\] so vertex is \[(\frac{11}{2},-\frac{9}{4})\]

OpenStudy (anonymous):

No , the options are (-7,-4) and (.5.5,2.25)